OpenStudy (angelina__nievse):

If x power of 2 = 40 what is the value of x

OpenStudy (mathstudent55):

\(x^2 = 40\) Solve for x. Is that the question?

OpenStudy (angelina__nievse):

Yes

OpenStudy (thomas123):

take the square root of 40.?

OpenStudy (mathstudent55):

An equation of the form \(x^2 = k\) where k is a number has the solutions \(x = \pm\sqrt k\)

OpenStudy (mathstudent55):

Take the positive and negative square roots of 40.

OpenStudy (angelina__nievse):

Ok thanks would it be 80

OpenStudy (mathstudent55):

You may need to simplify \(\sqrt {40}\) since there is a perfect square factor of 40.

OpenStudy (mathstudent55):

No. \(\sqrt {40} \) is the number that when you square you get 40. What number do you multiply by itself (square) to get 40? Let's see. 6 * 6 = 36 too low 7 * 7 = 49 too high \(\sqrt {40} \) is a number between 6 and 7.

OpenStudy (mathstudent55):

If you need an exact solution, then it is \(\sqrt {40} = \sqrt {4 \times 10} = 2 \sqrt{10}\) If you need an approximation, then \(\sqrt {40} \approx 6.324\) Either way, remember that for your question, you need both the positive and negative square roots.

OpenStudy (angelina__nievse):

Thanks you really helped

OpenStudy (mathstudent55):

You're welcome.